OpenStudy (anonymous):

Multiply both sides of the given equation by a number that will eliminate decimals or fractions. Then solve.

OpenStudy (anonymous):

\[0.3x ^{2} + 0.2x - 0.1=0\]

OpenStudy (anonymous):

That requires finding the least common multiple.

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

okay is that 10?

OpenStudy (anonymous):

10 will work.

OpenStudy (anonymous):

so I get \[3x^{2} + 2x - 11\]

OpenStudy (anonymous):

What is \(0.1 \times 10\)?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

so i get \[3x^{2} + 2x - 1=10\]

OpenStudy (anonymous):

Then what?

OpenStudy (mathmale):

Good. Now consolidate all terms on the left side of this equation. Try factoring. If that proves to be difficult, try the quadratic formula to identify the roots and then write the factors.

OpenStudy (anonymous):

The furthest I got was then.. \[3x^{2} + 2x - 11=0\]

OpenStudy (anonymous):

and i couldnt factor any further... no numbers will fit ... (3x,?)(x,?)

OpenStudy (mathmale):

Whoa. Sorry. Multiplying your original equation, 0.3x2+0.2x−0.1=0, by 10, will result in 3x^2 + 2x -1 = 0, not 3x^2 + 2x -11 = 0 Please go back and verify this. So, we're attempting to solve 3x^2 + 2x -1 = 0. I'd suggest factoring. The quadratic formula would be an alternative tool for factoring.

OpenStudy (anonymous):

don't we have to multiply 0 by ten? It says to multiply both sides of the equation by a number that will eliminate decimals or fractions...

OpenStudy (anonymous):

wait nevermind... 0 X 10= 0... not 10 lol.

OpenStudy (mathmale):

Again, try solving 3x^2 + 2x -1 = 0. I'd recommend factoring.